a very old man married a young women. Together, their ages add up to a century. The womans age is 17 years less than half the mans age. How old is each person?
Answer by Alan3354(69443)

a very old man married a young women. Together, their ages add up to a century. The womans age is 17 years less than half the mans age. How old is each person?
-------------
w + m = 100
w = m/2 - 17
-------------
m/2-17 + m = 100
3m/2 = 117
m = 78
w = 22